OpenStudy (anonymous):

Challenge: f(x,y)=x^(2)-2y^(2) Find a point (A,B) such that the plane tangent to z=f(x,y) at (A,B,f(A,B)) has an equation 3x+4y+5z=____. Fill in blank.

OpenStudy (anonymous):

\[f'(x,y)=2x-4y=0,x=2y\] \[(A,2A,A^2-2(2A)^2)\rightarrow(A,2A,-7A^2)\]

OpenStudy (anonymous):

\[f'(A,B)=2A-4(2A)=-6A\]

OpenStudy (anonymous):

\[3x+4y+5z=3(A)+4(2A)+5(-7A^2)=11A-35A^2\]

OpenStudy (anonymous):

tried

OpenStudy (anonymous):

where did the -7A come from?

OpenStudy (anonymous):

in the first step?

OpenStudy (anonymous):

nevermind i see now

OpenStudy (anonymous):

@UnkleRhaukus ,@cwrw238 ,@Callisto

OpenStudy (anonymous):

@Callisto ,@cwrw238

OpenStudy (anonymous):